### v1

# Digital quantum simulation of squeezed states via enhanced bosonic encoding in a superconducting quantum processor

Abstract
We present a fully digital approach for simulating single-mode squeezed states on a superconducting quantum processor using an enhanced bosonic encoding strategy. By mapping up to $2^{n}$ photonic Fock states onto $n$ qubits, our framework leverages Gray-code-based encodings to reduce gate overhead compared to conventional one-hot or binary mappings. We further optimize resource usage by restricting the simulation on Fock states with even number of photons only, effectively doubling the range of photon numbers that can be represented for a given number of qubits. To overcome noise and finite coherence in current hardware, we employ a variational quantum simulation protocol, which adapts shallow, parameterized circuits through iterative optimization. Implemented on the Zuchongzhi-2 superconducting platform, our method demonstrates squeezed-state dynamics across a parameter sweep from vacuum state preparation ($r=0$) to squeezing levels exceeding the Fock space truncation limit ($r>1.63$). Experimental results, corroborated by quantum state tomography and Wigner-function analysis, confirm high-fidelity state preparation and demonstrate the potential of Gray-code-inspired techniques for realizing continuous-variable physics on near-term, qubit-based quantum processors.
I Introduction
With continued development of quantum circuit-based architectures [1], these systems are poised to complement specialized quantum machines in addressing applications ranging from materials simulation [2, 3] to cryptographic protocols [4, 5] and pharmaceutical research [6]. In pursuit of a universal quantum processor, a variety of hardware platforms—ranging from superconducting circuits [7, 8] and trapped ions [9, 10] to neutral atoms [11, 12] and photonic systems [13, 14]—are under intensive development. Each implements quantum logic within the circuit model, whereby qubits undergo sequences of well-controlled gate operations. Superconducting qubits, for instance, have made rapid strides in qubit count and fidelity, aided by improvements in fabrication and control electronics. Trapped-ion platforms likewise provide high qubit connectivity and long coherence times, while neutral-atom arrays offer scalability through optical trapping techniques.
Intriguingly, photonic quantum computing has evolved along two main pathways: the discrete-variable (DV) approach [15] and the continuous-variable (CV) approach [16]. DV photonics treats single photons as qubits for a circuit-model calculation, but suffers from non-deterministic entangling gates [13]. In response, measurement-based quantum computing (MBQC) [14] protocols were devised, leveraging percolation [17] techniques to build large cluster states that enable universal DV quantum logic. Conversely, CV photonics encodes information in the infinite-dimensional Hilbert space of electromagnetic modes—often referred to as the “phase space.” This method also supports universal quantum computation, for example through the Gottesman–Kitaev–Preskill (GKP) encoding [18], which can implement a circuit model using CV states. Given these capabilities, it is both natural and beneficial to explore the inverse scenario—namely, simulating CV photonic (or more broadly, bosonic) physics on digital, qubit-based hardware. By mapping continuous-variable systems onto qubit registers, one can harness existing circuit-model devices to investigate and emulate complex quantum processes characteristic of photonic or vibrational modes.
To simulate CV systems on qubit-based quantum hardware, photonic Fock states must be explicitly mapped to multi-qubit states via encoding. A commonly considered direct method is one-hot (unary) encoding [19], where each Fock state $|n\rangle_{F}$ corresponds to a computational basis state $|0...010...0\rangle$. While intuitive, this approach incurs linear resource overhead: representing $d$ photon levels requires $d+1$ qubits, making it impractical for large $d$. To mitigate this, the Gray code (also known as the reflected binary code) has been studied [20]. Unlike one-hot encoding, which uses only $n$ of the available $2^{n}$ basis states, Gray codes fully exploit the Hilbert space, reducing qubit counts. Studies comparing various encoding methods [21, 22] reveal that the optimal choice depends critically on the target application. Recent work has also identified novel approaches such as encoding bosons via fermions followed by mapping fermionic states to qubits through the Jordan-Wigner transformation [23], though this method introduces trade-offs in resource allocation and gate complexity.
In this work, we present a generalized encoding framework that maps $2^{n}$ photonic states onto $n$ qubits. We analyze the efficiency of these encodings and demonstrate that the Gray code—a member of this encoding family—achieves high efficiency in bosonic system simulations, for which we provide theoretical justification. Furthermore, we implement a quantum simulation of a bosonic system on real hardware. Building on foundational work demonstrating two-photon squeezed state simulations with one-hot encoding [24], our protocol extends the accessible photon number to $6$ through optimized bosonic mapping. This advancement enables exploration of squeezing parameters up to $r=2$, operating beyond the regime where the truncated Fock space approximation breaks down due to saturation of maximum photon state population. To address the challenges of noise and limited coherence in quantum hardware, we implement a variational quantum simulation (VQS) protocol [25, 26, 27]. This approach leverages parameterized quantum circuits optimized through hybrid quantum-classical feedback, adaptively balancing circuit depth and simulation accuracy. By avoiding the strict gate sequence requirements of Suzuki-Trotter decomposition [28], the VQS framework enables robust evolution of squeezed states under realistic device.
The remainder of this paper is organized as follows. In Section II, we introduce our generalized encoding framework for mapping up to $2^{n}$ photonic Fock states onto $n$ qubits. Section II.1 outlines the construction of multi-qubit representations of the bosonic ladder operators and explains how to realize these encodings in practice. Section II.2 then discusses why Gray-code-based encodings are especially efficient for most bosonic simulations. In Section III, we demonstrate our method by simulating single-mode squeezing on a superconducting quantum processor (Zuchongzhi-2) [29] provided by QuantumCTek Co., Ltd. Section III.1 describes a specialized variant of the Gray code that encodes only even-photon Fock states, effectively doubling the highest photon number that can be represented. Section III.2 presents our variational quantum simulation (VQS) procedure, designed to reduce circuit depth while maintaining accuracy. In Section IV, we compare the simulated results to exact theoretical benchmarks and analyze the hardware performance. Finally, Section V summarizes our main conclusions.
II Encoding bosonic operator
A general bosonic Hamiltonian is expressed as $H=H\left(\{b_{i}^{\dagger},b_{j}\}\right)$, where $b_{i}$ ($b_{i}^{\dagger}$) is the annihilation (creation) operator for the $i$-th bosonic mode. These operators satisfy the commutation relation $[b_{i},b_{j}^{\dagger}]=\delta_{ij},$ where $\delta_{ij}$ is the Kronecker delta function. The entire Hilbert space of the system, $\mathbb{H}$ , which contains $m$ modes, is a direct product of the subspaces of each mode, $\mathcal{H}_{i}$, and is given by
$\mathcal{\mathbb{H}}$ $=\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes...\mathcal{H}_{m}.$ (1)
Our objective is to encode the single-mode space $\mathcal{H}$ onto a circuit model. The extension to multiple modes is straightforward by employing a larger quantum circuit with additional qubits. The single-mode Hilbert space $\mathcal{H}$ is spanned by the Fock basis $\{|i\rangle_{F}\}$, where $i$ are non-negative integers, and the basis state $|i\rangle_{F}$ represents a system containing $i$ bosons (e.g., photons). The subscript "$F$" distinguishes the Fock states from the computational basis (denoted without subscripts) used to encode the Fock states. For practical simulations, a cutoff is introduced at a maximum boson number, $N_{\text{Max}}$, such that $|N_{\text{Max}}\rangle$ is the highest allowed state, satisfying $b^{\dagger}|N_{\text{Max}}\rangle_{F}=0$.
II.1 General encoding
In this paper, we focus on encoding $N=2^{n}$ Fock states using $n$ qubits. Specifically, we encode a Fock space of dimension $N$ by mapping each Fock state to a unique $n$-qubit basis state. A code $c=\{c_{0},c_{1},...,c_{N-1}\}$ represents a permutation of the natural sequence $(0,1,...,N-1)$, resulting in $N!$ possible encoding schemes, including the natural encoding (also known as the binary code [21]).
To systematically organize these encoding schemes, we arrange all possible codes in dictionary order, forming a set $\mathcal{C}_{n}$. In this ordering, sequences are compared by examining each position sequentially; the sequence with the smaller value at the first differing position is considered smaller. Thus, the set is denoted as: $\mathcal{C}_{n}=\{\mathcal{C}_{n}^{[0]},\mathcal{C}_{n}^{[1]},...,\mathcal{C}_{n}^{[N!-1]}\}$ where $\mathcal{C}_{n}^{0}$ corresponds to the natural (binary) encoding.
Any code $c\in\mathcal{C}_{n}$ can be used to encode a single bosonic mode by representing the Fock state $|i\rangle_{F}$ with the computational basis state $|c_{i}\rangle$. The annihilation operator is encoded as:
$b$ $=\sum_{i=1}^{N-1}\sqrt{i}\mathcal{X}_{i,i-1}\mathcal{P}_{i},$ (2)
where $\mathcal{X}_{i,i-1}$ is a tensor product of Pauli $X$ operators acting on the qubits where $c_{i}$ and $c_{i-1}$ differ, and $\mathcal{P}_{i}=|c_{i}\rangle\langle c_{i}|$ is a projector onto the state $|c_{i}\rangle$ and given by
$\mathcal{P}_{i}$ $=\frac{1}{2^{n}}\sum_{\alpha=0}^{N-1}(-1)^{W(i\&\alpha)}\otimes_{k=0}^{n-1}Z_{k}^{\alpha_{k}},$ (3)
where $\alpha_{k}$ is the bit value on the $k$-th position of the integer $\alpha$, $W(*)$ denotes the Hamming weight, i.e., the number of “1”s in the binary representation, and “&\&&” represents the bitwise AND operation. In this paper, $X_{i}$, $Y_{i}$, and $Z_{i}$ denote the Pauli $X$, $Y$, and $Z$ operators acting on the $i$th qubit. The detailed derivation of Eq. 2 and Eq. 3 is provided in Appendix A.
Here, we provide an example of the encoding using $c=\mathcal{C}_{2}^{[1]}=\{0,1,3,2\}$. In this case, the working basis $|3\rangle$ (which corresponds to $|11\rangle$ in binary representation) represents the two-boson state $|2\rangle_{F}$, and $|2\rangle$ (which corresponds to $|10\rangle$ in binary representation) represents the three-boson state $|3\rangle_{F}$. The annihilation operator is given by:
$b$ $=|0\rangle\langle 1|+\sqrt{2}|1\rangle\langle 3|+\sqrt{3}|3\rangle\langle 2|.$ (4)
For calculating the middle term of Eq. 4, we have: $\mathcal{X}_{2,1}=X_{1}\otimes I_{0}$ since $1$ and $3$ differ at the second qubit (adopting 0-based indexing, the second index corresponds to $X_{1}$). The projector is given by $\mathcal{P}_{1}=\frac{1}{4}\left(\boldsymbol{1}-Z_{0}-Z_{1}+Z_{1}\otimes Z_{0}\right).$ The final expression of the encoded operator $b$ is:
$b$ $=\frac{1+\sqrt{3}}{4}X_{0}+i\frac{1-\sqrt{3}}{4}Y_{0}+\frac{\sqrt{2}}{4}X_{1}+\frac{\sqrt{2}}{4}iY_{1}$
$+\frac{1-\sqrt{3}}{4}Z_{1}X_{0}+i\frac{1+\sqrt{3}}{4}Z_{1}Y_{0}$
$-\frac{\sqrt{2}}{4}X_{1}Z_{0}-\frac{\sqrt{2}}{4}iY_{1}Z_{0}$
The expression of the encoded Hamiltonian is obtained by substituting the encoded $b$ into the formula $H=H(b^{\dagger},b)$. To reduce the number of terms in the final Hamiltonian expression, the key idea is to minimize the number of terms in $b$. Let us examine Eq. 2 and Eq. 3 more carefully to achieve this reduction.
As shown in Eq. 3, the number of terms in the projector $\mathcal{P}_{i}$ is fixed at $N$. All terms (indexed by different values of $\alpha$) in $\mathcal{P}_{i}$ are direct products of $Z$ and $I$ operators only. Moreover, for different $i$, the projectors $\mathcal{P}_{i}$ have identical forms except for different multiplicative factors. In fact, we have: $\sum_{i}\mathcal{P}_{i}=\boldsymbol{1}.$
Next, we discuss the effect of the term $\mathcal{X}_{i,i-1}$. The operator $b$ represents the "ladder" operation that destroys a boson in the system state. In the first quantization representation, it is a summation of all possible transitions between photon states that differ by one photon. Therefore, $\mathcal{X}_{i,i-1}$ must include at least one $X$ operator on a qubit where its product with $Z$ ($I$) (within the projector $\mathcal{P}_{i}$) generates $Y$ ($X$) terms. As a result, each term in $\mathcal{X}_{i,i-1}\mathcal{P}_{i}$ for different $i$ has a distinct form. However, it is still possible to combine some terms. If there is only one $X$ in each term of $\mathcal{X}_{i,i-1}$, there are only $\left(\begin{array}[]{c}n\\ 1\end{array}\right)=n$ different $\mathcal{X}_{i,i-1}$. Therefore, there are at most $nN=n2^{n}$ terms in $b$. If $\mathcal{X}_{i,i-1}$ contains more $X$ operators, the number of terms in Eq. 2 increases. For the case of $n=3$, we counted the number of terms in $b$ for all codes in $\mathcal{C}_{3}$ and presented the results in Table 1. Out of the total $N!=40320$ codes, the encoded $b$ contains the minimal number of terms $nN=24$ in $4032$ codes. We define a subset $\mathcal{D}_{n}\subset\mathcal{C}_{n}$ that contains codes where there is only one $X$ in each $\mathcal{X}_{i,i-1}$ in Eq. 2. The Hamming distance between neighboring numbers in the codes in $\mathcal{D}_{n}$ is 1. We counted that there are 144 entries in $\mathcal{D}_{3}$.
Number of terms in $b$ Counts codes
24 4032
32 14784
40 14784
48 6720
Table 1: Distribution of the number of terms in the encoded annihilation operator $b$ across different encoding schemes in $\mathcal{C}_{3}$.
II.2 Gray code
In many previous works, the Gray code is often explained simply as a code where the Hamming distance between neighboring numbers is $1$, which is a definition captured by $\mathcal{D}_{n}$ in this paper. However, here we clarify that the Gray code is just one specific instance within $\mathcal{D}_{n}$. The Gray code is defined by the recursive generation formula given by [20]
$G_{n}$ $=(\boldsymbol{0}\cdot G_{n-1},\boldsymbol{1}\cdot\bar{G}{}_{n-1}),$ (5)
with $G_{1}=(0,1)$, where $\bar{G}_{n-1}$ represents the reverse order of $G_{n-1}$. The advantage of using the Gray code $G_{n}$ lies in the established rules for generating a specific instance within the $\mathcal{D}_{n}$ codes. As discussed above, $\mathcal{D}_{n}$ is an excellent option for encoding the operator $b$. These codes are prime candidates for studying Hamiltonians, such as those involving an external field $g(b+b^{\dagger})$ or the generator of the displacement operator $\alpha b^{\dagger}-\alpha^{*}b$.
If the Hamiltonian container higher order terms
$b^{k}$ $=\sum_{i=0}^{N-1-k}\sqrt{\frac{(i+k)!}{i!}}|i\rangle_{F}\langle i+k|_{F}\text{, }(k\leq N-1),$ (6)
we have different strategies. As shown in Eq. 6, in the summation terms of the expression $b^{k}$, the $i$-th term is coupled with all terms $i+k$, $i+2k$,…, and thus the entire summation can be partitioned into $k$ distinct groups. Each group contains terms that are separated by a multiple of $k$. We rewritten the expression of $b^{k}$ in a form of summation of different groups as
$b^{k}$ $=\sum_{\delta=0}^{k-1}\left[\sum_{i=0}^{I(k,\delta)}\sqrt{\frac{\left(ik+k+\delta\right)!}{(ik+\delta)!}}|ik+\delta\rangle_{F}\langle(i+1)k+\delta|_{F}\right],$ (7)
where $I(k,\delta)=\left\lfloor\frac{1}{k}(N-1-\delta)-1\right\rfloor$. We define a code to be $k$-fold if, within the encoded sequence, any two codewords that are separated by $k$ positions have a Hamming distance of $1$. A $k$-fold encoding is characterized by partitioning the code into $k$ groups, where each group consists of elements that are spaced $k$ apart. This grouping ensures that the encoding operator adheres to the structure defined in Eq. 7.
As an example, when $k=2$, the summation is divided into two groups: one representing the sum of the odd-indexed terms and the other representing the sum of the even-indexed terms, as illustrated by
$b^{2}$ $=\underbrace{\left(\sqrt{6}|1\rangle_{F}\langle 3|_{F}+\sqrt{20}|3\rangle_{F}\langle 5|_{F}+...\right)}_{\text{odd terms}}$
$+\underbrace{\left(\sqrt{2}|0\rangle_{F}\langle 2|_{F}+\sqrt{12}|2\rangle_{F}\langle 4|_{F}+...\right)}_{\text{even terms}}.$ (8)
In this scenario, a 2-fold code is naturally suited for encoding $b^{2}$, as anticipated. As illustrated in Fig. 1, the number of terms in $b^{2}$ varies with the number of qubits $n$ across several representative codes. Notably, the binary code $\mathcal{C}_{n}^{[0]}$ is intrinsically a 2-fold code and consistently exhibits the lowest cost among the compared codes. In contrast, $D_{n}^{[0]}$, which is the first code in dictionary order that ensures a Hamming distance of 1 between adjacent numbers, incurs a higher cost, resulting in a greater number of terms. $C_{n}^{[5]}$ is included for comparison. As shown in the figure, the Gray code $G_{n}$ also performs favorably, maintaining a relatively low number of terms.
Figure 1: Number of terms in the encoding operator $b^{2}$ as a function of qubit number $n$. The lines represent different codes: $\mathcal{C}_{n}^{[0]}$ (blue, solid with circles), $\mathcal{C}_{n}^{[5]}$ (orange, solid with circles), $G_{n}$ (green, dashed with crosses), and $D_{n}^{[0]}$ (red, dashed with circles) which is the first code in dictionary order that ensures a Hamming distance of 1 between adjacent numbers.
However, one issue arises: why does the Gray code $G_{n}$ still perform well, given that $G_{n}\in\mathcal{D}_{n}$ and we believe that 1-fold codes are not suitable for encoding $b^{2}$? We state that $G_{n}$ is a special code that is also effective for encoding $b^{2}$. To clarify this, we present an example for the case where $n=3$. The binary representations of the numbers in $G_{3}$ are regrouped into even-indexed and odd-indexed terms and are listed in Table 2. As can be seen, apart from the rightmost qubit, the remaining two qubits (highlighted in red) still form the Gray code $G_{2}$. In this configuration, there is a fixed operator $X$ acting on the 0-th bit position. We will soon see that this additional $X$ does not adversely affect the encoding process.
even terms odd terms
000 001
011 010
110 111
101 100
Table 2: Binary representations of $G_{3}$ divided into even and odd terms.
According to Eq. 2, the operator $b^{2}$ in this encoding ($G_{n}$) can be expressed as:
$b^{2}$ $=\sum_{i=1}^{N-1}\sqrt{i}\mathcal{X}_{i,i-1}^{\prime}X_{0}\mathcal{P}_{i}X_{0}\sum_{j=1}^{N-1}\sqrt{j}\mathcal{X}_{j,j-1}^{\prime}\mathcal{P}_{j}$
$=\left(\sum_{i=1}^{N-1}(-1)^{i_{0}}\sqrt{i}\mathcal{X}_{i,i-1}^{\prime}\mathcal{P}_{i}\right)\left(\sum_{j=1}^{N-1}\sqrt{j}\mathcal{X}_{j,j-1}^{\prime}\mathcal{P}_{j}\right)$ (9)
where $\mathcal{X}_{i,i-1}^{\prime}$ represents the collection excluding the 00-th qubit position, and $i_{0}$denotes the value of the 0-th bit in the binary representation of $i$. The second line in Eq. 9 is derived by utilizing the identity $X_{0}Z_{0}^{\delta}X_{0}=(-1)^{\delta}Z_{0}^{\delta}$ for any $\delta$. This expression demonstrates that if there is an $X$ operator acting on each term of $\mathcal{X}_{i,i-1}$ at a given bit position, the number of terms in the operator’s expression remains unchanged. In Appendix B, we shown that $G_{n}$ can be used as any $2^{\xi}$-fold code, therefor $G_{n}$ is suit for encoding the operator $b^{l}$ with $l=2^{\xi}$.
Next, we present a counterexample to examine the limitations of $G_{n}$. Specifically, when $k=3$, the local symmetry of $G_{n}$ is lost, and a 3-fold code should be the most suitable for encoding in this scenario. As shown in Fig. 2, we display the number of terms in $b^{3}$ as a function of the number of qubits $n$. It is evident that for $n\geq 3$, $G_{n}$ is no longer the optimal encoding method.
It is important to note that the figure lacks data for the 3-fold code at $n=6$ and $n=8$. This omission is due to the absence of corresponding 3-fold codes for these values of $n$. Efficiently finding 3-fold codes is mathematically challenging, and thus, we do not discuss this in further detail here.
Figure 2: Number of terms in the encoding operator $b^{3}$ as a function of the number of qubits $n$. The cyna dashed line represents the Gray code $G_{n}$, and the orange dashed line represents the 3-fold code.
III Squeezed state simulation
III.1 Encoded Hamiltonian
For practical quantum simulation, we consider encoding the squeezed state dynamics using a Gray code method. In previous studies [24], a one-hot encoding scheme was used to represent a single-mode squeezed state. In a one-hot scheme with $n$ qubits, each basis state features exactly one qubit in the $|1\rangle$ state and the remainder in $|0\rangle$, allowing the circuit to represent photonic states with up to $n-1$ photons (yielding $n$ distinct states). In contrast, Gray code encoding exploits the full $2^{n}$ computational basis, enabling simulation of photonic states with up to $12^{n}-$ 1 photons.
The single-mode squeezed state can be generated by applying a squeezing operator $S(z)$ to the vacuum state. The squeezing operator is defined as
$S(z)$ $=\exp\left[\frac{1}{2}\left(z^{*}b^{2}-zb^{\dagger 2}\right)\right],$ (10)
where $z=re^{i\varphi_{z}}$ is the complex squeezing parameter (with $r=|z|$ characterizing the squeezing level and complex angle $\varphi_{z}$ determine the squeezing direction in the phase space). Here $b$ is the photon annihilation operator for the mode. The squeezing process can be viewed as an evolution under a Hamiltonian $H$: specifically $S=e^{-iHt}|_{t=r},$ evaluated at time $t=r$. The Hamiltonian that generates this transformation is given by
$H$ $=\frac{1}{2}\left[e^{-i\left(\varphi_{z}-\frac{\pi}{2}\right)}b^{2}-e^{i\left(\varphi_{z}+\frac{\pi}{2}\right)}b^{\dagger 2}\right].$ (11)
Applying the operator $S(z)$ to the vacuum $|0\rangle$ (the single-mode Fock vacuum) yields an analytical expression for the squeezed state [30], and it is given by
$|z\rangle$ $=\frac{1}{\sqrt{\mu}}\exp\left(-\frac{\nu}{2\mu}b^{\dagger 2}\right)|0\rangle,$ (12)
where $\mu=\cosh r$ and $\nu=e^{i\varphi_{z}}\sinh r$. Notably, the Fock basis expansion of $|z\rangle$ contains only even-photon-number components. Exploiting this property, we improve our encoding efficiency by representing only even-photon Fock states. Consequently, the $2^{n}$ available codewords can encode Fock states with photon numbers up to $2(2^{n}-1)$; in other words, with $n$ qubits, we can represent photon numbers as high as twice the usual limit.
For example, with $n=2$ qubits, a one-hot encoding can represent up to $1$ photon, while a Gray code can normally represent up to $3$ photons. If we restrict to even photon numbers, those two qubits can instead represent the four Fock states $\{|0\rangle_{F},|2\rangle_{F},|4\rangle_{F},|6\rangle_{F}\}$ (0, 2, 4, and 6 photons). We assign these Fock states to the 2-qubit computational basis states using the Gray code sequence $G_{2}=\{|00\rangle,|01\rangle,|11\rangle,|10\rangle\}$. According to the general encoding formula Eq. 9, the annihilation operator $b^{2}$ under this encoding is expressed as:
$b^{2}$ $=\frac{\sqrt{2}+\sqrt{30}}{4}X_{0}+\frac{\sqrt{12}}{4}X_{1}+\frac{\sqrt{2}-\sqrt{30}}{4}iY_{0}+\frac{\sqrt{12}}{4}iY_{1}$
$+\frac{\sqrt{2}-\sqrt{30}}{4}Z_{1}X_{0}+\frac{\sqrt{2}+\sqrt{30}}{4}iZ_{1}Y_{0}$
$-\frac{\sqrt{12}}{4}X_{1}Z_{0}-\frac{\sqrt{12}}{4}iY_{1}Z_{0}.$ (13)
Substituting this encoded form of $b^{2}$ into the Hamiltonian $H$ in Eq. 11, we obtain the qubit-encoded Hamiltonian for the squeezed state, and it is given by: $H=H_{R}+H_{I},$ where
$H_{R}$ $=\cos\varphi_{z}\left(\frac{\sqrt{30}-\sqrt{2}}{4}Y_{0}-\frac{\sqrt{3}}{2}Y_{1}\right.$ (14)
$\left.-\frac{\sqrt{30}+\sqrt{2}}{4}Z_{1}Y_{0}+\frac{\sqrt{3}}{2}Y_{1}Z_{0}\right)$
and
$H_{I}$ $=\sin\varphi_{z}\left(\frac{\sqrt{30}+\sqrt{2}}{4}X_{0}+\frac{\sqrt{3}}{2}X_{1}\right.$ (15)
$\left.-\frac{\sqrt{30}-\sqrt{2}}{4}Z_{1}X_{0}-\frac{\sqrt{3}}{2}X_{1}Z_{0}\right).$
Thus, we have explicitly encoded the squeezing Hamiltonian in terms of two-qubit operations. This encoded Hamiltonian can now be used in the circuit model to simulate the squeezed state dynamics on a quantum processor.
III.2 Variational quantum simulation
For near-term quantum devices where noise and limited gate depth hinder deep circuit implementations, the traditional Suzuki–Trotter method becomes impractical. Instead, we adopt Variational Quantum Simulation (VQS) [25, 26, 27], which uses a shallow, parameterized circuit to approximate time evolution. We provide here a concise introduction to the principle of VQS.
III.2.1 Ansatz for evolution
Once the physical system is encoded into a circuit model, the Hamiltonian can be expressed as a weighted sum of Pauli strings,
$H$ $=\sum_{i}\xi_{i}P_{i},$ (16)
where each $\xi_{i}$ is a real coefficient and each operator $P_{i}$ is given by a tensor product of Pauli matrices (i.e. $P_{i}=\sigma_{n-1}\otimes...\otimes\sigma_{0}$ with $\sigma_{i}\in\{I,X,Y,Z\}$ acting on the corresponding qubit in the circuit). Given an initial state $|\psi_{0}\rangle$, the exact time-evolved state under $H$ is
$|\psi_{t}\rangle$ $=e^{-iHt}|\psi_{0}\rangle.$
In the fault-tolerant era, one standard approach is the Suzuki–Trotter decomposition:
$e^{-iHt}$ $\approx\left(\prod_{j}e^{-iP_{j}\xi_{j}\frac{t}{N_{\tau}}}\right)^{N_{\tau}},$
with large $N_{\tau}$. However, on near-term quantum devices subject to noise and gate imperfections, large circuit depths quickly degrade the fidelity of the final state. A variational strategy addresses this by approximating the time evolution via a shallower, parametrized circuit. This method is often referred to as VQS [25, 26, 27]—an approach related to the philosophy of variational quantum eigensolvers (VQE) [31, 32].
In a VQS approach, one replaces the exact propagator $e^{-iHt}$ by a parametrized, shallower-depth ansatz circuit
$\Gamma(\boldsymbol{\theta})$ $=\Gamma(\theta_{0},\theta_{1},...,\theta_{N-1}),$ (17)
where each $\theta_{i}=\theta_{i}(t)$ is a time-dependent variational parameter. The total number of these parameters is kept deliberately small to maintain low-depth circuits. The goal is then to evolve the system in time by adjusting $\boldsymbol{\theta}(t)$ so that
$|\psi_{t}\rangle$ $\approx|\psi(\boldsymbol{\theta})\rangle=\Gamma(\boldsymbol{\theta})|\psi_{0}\rangle.$ (18)
III.2.2 Equations of motion (EOM) of VQS
This strategy aims to capture the essential dynamics of the system without incurring the large circuit depths required by direct Trotter decomposition. The time-dependent variational principle which derives the Schrödinger equation [25] is given by
[…]
where the $L=\langle\psi(\boldsymbol{\theta})|i\frac{\partial}{\partial t}-H|\psi(\boldsymbol{\theta})\rangle$ is the Lagrangian of the system. With considering Eq. 18, the Lagrangian in this representation is given by
$L(\boldsymbol{\theta},\dot{\boldsymbol{\theta}})$ $=i\sum_{k}\langle\psi(\boldsymbol{\theta})|\frac{\partial|\psi(\boldsymbol{\theta})\rangle}{\partial\theta_{k}}\dot{\theta}_{k}-\langle\psi(\boldsymbol{\theta})|H|\psi(\boldsymbol{\theta})\rangle.$ (20)
Therefore the variational principle in Eq. 19 derives the equation set
$\sum_{q}M_{k,q}\dot{\theta}_{q}$ $=V_{k},$ (21)
where
$M_{kq}$ $=i\eta\frac{\partial\langle\psi|}{\partial\theta_{k}}\frac{\partial|\psi\rangle}{\partial\theta_{q}}+h.c.$ (22)
$V_{k}$ $=\eta\frac{\partial\langle\psi|}{\partial\theta_{k}}H|\psi\rangle+h.c.$ (23)
[…]
III.2.3 Circuit for calculation $M$ and $V$
We consider a layered or factorized ansatz of the form
$\Gamma(\theta)$ $=\Gamma_{N-1}(\theta_{N-1})\Gamma_{N-2}(\theta_{N-2})...\Gamma_{0}(\theta_{0}),$
where each $\Gamma_{i}(\theta_{i})=A_{i}R_{i}(\theta_{i})$ is composed of:
  1. 1.
    A fixed (time-independent) product of gates $A_{i}$
  2. 2.
    A single-qubit rotation $R_{i}(\theta_{i})=e^{-i\sigma_{i}\frac{\theta_{i}}{2}}$ about a specified axis $\sigma_{i}\in\{X,Y,Z\}$.
It follows that
$\frac{d}{d\theta_{i}}R_{i}(\theta_{i})$ $=\frac{1}{2i}R_{i}\sigma_{i}.$ (24)
From this, one can derive
$\frac{\partial\Gamma(\boldsymbol{\theta})}{\partial\theta_{i}}$ $=\Gamma_{N-1}...\Gamma_{i+1}\Gamma_{i}\sigma_{i}\Gamma_{i-1}...\Gamma_{0}.$ (25)
Using the above derivative rule in the definition of $M_{kq}$ (Eq. 22), we obtain
$M_{kq}$ $=\frac{i}{4}\langle\psi_{0}|\Gamma_{0}^{k-1{\dagger}}\sigma_{k}^{\dagger}\Gamma_{k}^{q-1\dagger}\sigma_{q}\Gamma_{k}^{q-1}\Gamma_{0}^{k-1}|\psi_{0}\rangle+h.c.,$ (26)
where we have used the notation $\Gamma_{k}^{k+l}=\Gamma_{k+l}...\Gamma_{k+1}\Gamma_{k}$. To measure $M_{kq}$ on a quantum processor, one can implement the circuit in Fig.  3, where an ancillary qubit is introduced and used to perform a measurement-based estimation of the above amplitude. Denoting the measured probabilities of the ancilla as $p_{0}$ and $p_{1}$ (for observing 00 and $1$ respectively), one has
$M_{kq}$ $=\frac{1}{2}(p_{0}-p_{1})=p_{0}-\frac{1}{2}.$ (27)
This procedure is repeated for all required pairs $(k,q)$ to construct the full matrix $\boldsymbol{M}$.
{quantikz}
[…]
Figure 3: The ancilla-based measurement circuit used to estimate each matrix element $M_{kq}$.
Similarly, one obtains from the definition of $V_{k}$ (Eq. 23) that $V_{k}=2\sum_{m}\xi_{m}V_{km},$ where
$V_{km}$ $=\frac{i}{4}\langle\psi_{0}|\Gamma_{0}^{k-1\dagger}\sigma_{k}\Gamma_{k}^{n-1\dagger}P_{m}\Gamma_{k}^{n-1}\Gamma_{0}^{k-1}|\psi_{0}\rangle+h.c..$ (28)
As with $M_{kq}$, the circuit for estimating $V_{km}$ is nearly the same, except that $P_{m}$ is inserted in place of $\sigma_{q}$. This is illustrated in Fig. 4.
{quantikz}
[…]
Figure 4: Circuit diagram for measuring the vector components $V_{km}$.
III.2.4 Variational Hamiltonian ansatz (VHA)
In the preceding discussions, we introduced the VQS framework and highlighted that one must specify the functional form of the variational ansatz $\Gamma(\boldsymbol{\theta})$. Choosing this ansatz is a nontrivial challenge, as naive or random circuit structures can lead to two major problems. First, a poorly chosen ansatz may yield trivial parameter dynamics, producing $\dot{\boldsymbol{\theta}}(t)=0$ for all $t$ and implying no meaningful evolution. Second, even when the EOM do evolve in time, a suboptimal ansatz might fail to capture key features of the quantum state, so that no matter how small the time-step $\Delta t$ we used in incrementing parameters, the resulting simulation deviates substantially from the true $e^{-iHt}|\psi_{0}\rangle$. These issues have been extensively discussed in Ref.  [27]. One robust strategy to mitigate them is the so-called Variational Hamiltonian Ansatz (VHA) [34, 35]. In VHA, we construct the parametrized circuit by mimicking short Trotter-like steps of the original Hamiltonian. Concretely, we consider a circuit with multiple layers (“depth” $n_{d}$) of exponentials of each Pauli string. That is, we write:
$\Gamma(\boldsymbol{\theta})$ $=\prod_{d=0}^{n_{d}-1}\left[\prod_{j}\exp\left(-iP_{j}\theta_{dj}\right)\right].$ (29)
Each layer thus contains exponentials of the same set of Pauli operators $\{P_{j}\}$ appearing in $H$. In effect, one can view each layer as a “Trotter slice,” except that we allow different, variationally optimized angles $\theta_{dj}$ per layer. This approach has several advantages. It is systematically improvable: by increasing $n_{d}$, one enlarges the variational space and can asymptotically recover exact dynamics. It also has physical interpretability: each parameter $\theta_{dj}$ directly corresponds to how strongly the relevant Hamiltonian term acts. Moreover, it reduces the risk of stagnation because the VQS solution coincides with the exact propagator for vanishing small times [27].
In practice, one sets the desired number of layers $n_{d}$ by starting with the smallest possible $(n_{d}=1)$ and incrementing until the approximation quality is acceptable. We adopt the VHA in our calculations.
III.2.5 Summary of the VQS procedure
The steps to implement VQS are concluded as follows:
  1. 1.
    Initialization: Set $\boldsymbol{\theta}(t=0)$ so that $\Gamma(\boldsymbol{\theta})=\boldsymbol{1}$ is effectively the identity. Therefore we have $|\psi_{t=0}\rangle=|\psi_{0}\rangle$.
  2. 2.
    Measurement of $\boldsymbol{M}$ and $\boldsymbol{V}$: For the current parameter set $\boldsymbol{\theta}(t)$, run the ancilla-aided circuits in Fig. 3 and Fig. 4 to obtain all elements $M_{kq}$ and $V_{k}$. Build the matrix $\boldsymbol{M}$ and vector $\boldsymbol{V}$ from these measurements.
  3. 3.
    Update rule: Solve the linear system $\boldsymbol{M}(t)\dot{\boldsymbol{\theta}}(t)=\boldsymbol{V}(t)$ for $\dot{\boldsymbol{\theta}}(t)$. In practice, one can use standard linear algebra methods.
  4. 4.
    Increment parameters: Propagate by a small time step $\Delta t$, i.e. $\boldsymbol{\theta}(t+\Delta t)\approx\boldsymbol{\theta}(t)+\dot{\boldsymbol{\theta}}(t)\Delta t$. After this, go to step 2 to start new calculation for the next time step.
  5. 5.
    Approximate the evolved state: At any intermediate time $t$, the circuit $\Gamma(\boldsymbol{\theta})$ approximates the true evolution $e^{-iHt}$. Thus $|\psi_{t}\rangle\approx\Gamma\left(\boldsymbol{\theta}(t)\right)|\psi_{0}\rangle$.
IV results
IV.1 Fidelity of VQS
Figure 5: Fidelity vs. $r=|z|$ for the truncated exact solution (blue dashed), VQS simulation (blue dashed), and quantum hardware Zuchongzhi-2 (green). The vertical red dashed line marks $r_{0}\approx 1.63$, where $n^{s}(r_{0})=6$.
We employ the Gray code encoding to simulate the evolution of a vacuum state under single-mode squeezing. As discussed before, the encoding is restricted to even-photon Fock states, thus omitting contributions from odd photon numbers. In the illustrative examples presented here, the truncated subspace spans four Fock states $\{|0\rangle_{F},|2\rangle_{F},|4\rangle_{F},|6\rangle_{F}\}$, allowing us to capture photon numbers up to six. For simplification, the squeezing parameter $z$ is taken to be purely imaginary ($z=it$), so that the real component of the squeezing Hamiltonian vanishes. Empirically, a single variational layer ($n_{d}=1$) suffices for accurately capturing the relevant dynamics.
To assess simulation accuracy, we define the fidelity of a state $|\psi\rangle$ (restricted to this four dimensional subspace) as $F(|\psi\rangle)=|\langle\psi|z\rangle|^{2}$, where $|z\rangle$ is the exact squeezed state (see Eq. 12), truncated to the same maximum photon number of six in this case.
Figure 5 plots the fidelity versus $r=|z|$. The curve labeled “Exact” represents the mathematical wavefunction of the squeezed state, forcibly truncated at six photons. Because a true single-mode squeezed state has an average photon number $n^{s}(z)=\sinh^{2}r$, once $n^{s}(z)$ grows beyond $6$, the truncated subspace no longer captures a significant portion of the total wavefunction. Consequently, this truncated “Exact” fidelity drops below unity for large $r$. A vertical red dashed line indicates $r_{0}\approx 1.63$, where $n^{s}(r_{0})=6$; beyond that point, truncating at six photons omits a substantial fraction of the state.
The VQS (blue dashed line) closely follows the truncated exact curve across nearly the entire range. Only near $r\approx 1.25$ does a minor discrepancy arise. Even then, the deviation remains small, underscoring that our Gray code encoding and single-layer variational approach can reliably capture squeezed-state evolution up to quite large $r$. For $r$ beyond the red dashed line, the average photon number exceeds six, causing the truncated state (and hence our benchmark) to lose fidelity more substantially.
IV.2 Quantum state tomography
Figure 6: Reconstructed two-qubit density matrix $\rho$ at $r=0.5$. Red bars represent the experimentally measured data (via quantum state tomography on the Zuchongzhi-2 processor), whereas gray bars show the corresponding truncated theory values.
[…]
$\rho$ $=\sum_{ij}\frac{\text{tr}\left(\rho\sigma_{i}^{1}\otimes\sigma_{j}^{0}\right)}{4}\sigma_{i}^{1}\otimes\sigma_{j}^{0},$ (30)
where each coefficient $\text{tr}\left(\rho\sigma_{i}^{1}\otimes\sigma_{j}^{0}\right)$ is determined via measurements in the corresponding Pauli basis. Although the total number of measurement settings scales exponentially with qubit count, tomography remains tractable for two qubits. For larger systems, an ancilla-based Hadamard test [36] may be more efficient, but it requires extra qubits and additional entangling gates, which can introduce further noise.
In our experiment, we implement direct measurements of each two-qubit Pauli basis on the Zuchongzhi-2 device, collecting 50,000 shots per basis to bolster the statistical reliability. We further optimize performance by transpiling the parameterized circuit (with the calibrated parameters $\boldsymbol{\theta}$) into Zuchongzhi-2’s native gate set $\{X2P,X2M,Y2P,Y2M,R_{z},CZ\}$ using bqskit [37], thus minimizing gate layers and reducing error accumulation. In Appendix C, we will give an example of this calculation.
Figure 6 displays the reconstructed two-qubit density matrix $\rho$ for $r=0.5$. Although certain components, such as $Z\otimes I$ and $Y\otimes Z$, deviate noticeably from their ideal values (e.g., in amplitude and sign), the overall fidelity $\text{tr}\left(\rho|z\rangle\langle z|\right)$ remains above $0.9$. This indicates that the core features of the squeezed state are captured relatively well. Referring back to Fig. 5 (green dot-dashed line), we see that across a broad range of $r$, the hardware fidelity remains high, agreeing well with the ideal truncated solution until the effective photon number grows too large for our 4-dimensional qubit encoding. These results confirm that, even with hardware imperfections and finite sampling, the proposed encoding and variational approach faithfully reproduce a significant portion of the squeezed state’s target distribution.
IV.3 Wigner spectra
Figure 7: Wigner distributions of the single-mode squeezed state at $|z|=0.5$. From left to right: (i) experimentally obtained result on Zuchongzhi-2, (ii) theoretical (exact) truncation to a six-photon Fock subspace, and (iii) the ideal infinite-dimensional state.
Once the two-qubit density matrix has been reconstructed via QST, we can evaluate the corresponding Wigner-function [38, 39] to visualize the system’s phase-space characteristics. Figure 7 illustrates the Wigner distribution at $r=0.5$ for three cases (from the left to the right in the figure): (i) the experimentally obtained state on the QPU Zuchongzhi-2, (ii) the theoretical (exact) result truncated to a six-photon Fock subspace, and (iii) the exact infinite-dimensional squeezed state.
We can factor the single-mode squeezing operator as $S(z)=R(\frac{\varphi_{z}}{2})S(r)R(-\frac{\varphi_{z}}{2})$, where $S(r)$ compresses the $x$-quadrature (and stretches $p$). Since $\varphi_{z}=\frac{\pi}{2}$ here, the overall operation is rotated by $\frac{\pi}{4}$, leading to the 45° tilt in the Wigner plots, as shown in Fig. 7.
From the infinite-dimensional perspective, a squeezed state at this moderate squeezing value exhibits a purely positive Wigner-function. The finite truncation at six photons, however, introduces slight negative “ring-like” regions toward the fringes, reflecting the imperfect capture of higher photon-number components. These negative pockets do not represent genuine nonclassical phenomena at this moderate squeezing level; rather, they are an artifact of disallowing photon numbers above six.
Experimentally, our reconstructed Wigner-function largely retains the squeezed profile but shows more widespread negative dips than the truncated case. This discrepancy underscores the practical effects of both device noise and the same finite-photon cutoff. Nevertheless, the central squeezed peak and its diagonal orientation are readily discernible, indicating that the main features of the targeted squeezed state remain intact under our encoding and simulation procedure.
Overall, these Wigner spectra confirm that, despite hardware imperfections and photon-number truncation, the essential squeezed-state characteristics (notably, the elongated shape in phase space) are successfully reproduced. This consistency with theoretical expectations, even for a relatively modest number of computational basis states, further validates our bosonic encoding strategy and demonstrates the feasibility of exploring continuous-variable phenomena on digital quantum processors.
V conclusion
In this work, we have developed a comprehensive framework to encode $2^{n}$ bosonic modes using $n$ qubits. Within this paradigm, the Gray code stands out by achieving excellent efficiency—minimizing the number of Pauli strings needed to represent the operator $b^{l}$ (with $l=2^{\xi}$)—due to its inherent symmetry. This efficiency directly translates into reduced circuit depth and lower gate complexity.
Furthermore, we introduced a truncated encoding scheme that maps only even-photon Fock states, thereby effectively doubling the maximum photon number accessible per qubit. For instance, with $n=2$ qubits, our method can simulate photonic states with up to $6$ photons. This encoding strategy was integrated into a variational quantum simulation (VQS) framework, where shallow circuits and ancilla-assisted measurements were used to iteratively update the time-dependent variational parameters ($M_{kq}$ and $V_{km}$). The VQS results show good agreement with exact simulations, demonstrating the robustness of this approach.
Experimental implementations on the Zuchongzhi-2 processor further validated our method by producing high-fidelity squeezed states at moderate squeezing levels, as confirmed by both quantum state tomography and Wigner-function measurements. Overall, our results not only establish the feasibility of simulating continuous-variable (CV) states on digital, qubit-based quantum processors but also underscore the promising potential of Gray-code methods for a wide range of bosonic Hamiltonians.
Acknowledgements.
This work was supported by the Beijing Nova Program (Grants No. 20220484128 and 20240484652). Computational resources were provided by the Chaohumingyue Hefei QC-HPC Hybrid Computing Center.
Appendix A Projector
We begin with the expression for the annihilation operator in the Fock basis:
$b$ $=\sum_{i=1}^{N-1}\sqrt{i}|i-1\rangle_{F}\langle i|_{F}.$ (31)
Consider the encoding mapping $\{|c_{i}\rangle\}$. Under this encoding, the operator $|i-1\rangle_{F}\langle i|_{F}$ is mapped to:
$|i-1\rangle_{F}\langle i|_{F}$ $=|c_{i-1}\rangle\langle c_{i}|$
$=O_{n-1}\otimes O_{n-2}\otimes...\otimes O_{0},$ (32)
where $O_{\alpha}$ is an operator acting on the $\alpha$-th qubit. The explicit form of $O_{\alpha}$ depends on the values of $c_{i-1}$ and $c_{i}$ at the $\alpha$-th bit of their binary representations. Specifically, $O_{\alpha}$ is defined as [40]:
$O_{\alpha}=\left\{\begin{array}[]{c}|0\rangle\langle 0|=(\boldsymbol{1}+Z)/2\text{ if }c_{i-1}^{\alpha}=0\text{ and }c_{i}^{\alpha}=0\\ |1\rangle\langle 1|=(\boldsymbol{1}-Z)/2\text{ if }c_{i-1}^{\alpha}=1\text{ and }c_{i}^{\alpha}=1\\ |1\rangle\langle 0|=(X-iY)/2\text{ if }c_{i-1}^{\alpha}=1\text{ and }c_{i}^{\alpha}=0\\ |0\rangle\langle 1|=(X+iY)/2\text{ if }c_{i-1}^{\alpha}=0\text{ and }c_{i}^{\alpha}=1\end{array}\right..$ (33)
We can simplify the last two cases to obtain a more compact form. For example, the operator $|1\rangle\langle 0|$ can be rewritten as:
$|1\rangle\langle 0|$ $=\frac{1}{2}(X-ZX)$
$=\frac{1}{2}X(\boldsymbol{1}+Z)=X\cdot|0\rangle\langle 0|.$ (34)
This demonstrates that the transformation operator $|1\rangle\langle 0|$ can be expressed as a projection onto the $|0\rangle$ state followed by a bit flip $X$ on the relevant qubit. Similarly, we have: $|0\rangle\langle 1|=X\cdot|1\rangle\langle 1|.$ Therefore $|c_{i-1}\rangle\langle c_{i}|$ can be written as:
$|c_{i-1}\rangle\langle c_{i}|$ $=\mathcal{X}_{i-1,i}\mathcal{P}_{i},$ (35)
where $\mathcal{P}_{i}=|c_{i}\rangle\langle c_{i}|$ is the projector onto the state $|c_{i}\rangle,$ and $\mathcal{X}_{i-1,i}$ is defined as described in the main text. The state projector $\mathcal{P}_{i}$ can be further expressed in terms of Pauli $Z$ operators as:
$\mathcal{P}_{i}$ $=\frac{1}{2^{n}}\left[\boldsymbol{1}+(-1)^{i_{n-1}}Z\right]\otimes...\otimes\left[\boldsymbol{1}+(-1)^{i_{0}}Z\right],$ (36)
where $i_{\alpha}$ is the value of the $\alpha$-th bit (0-based) of the integer $i$ in its binary representation. Expanding the tensor products yields Eq. 3 from the main text.
Appendix B $2^{\xi}$-fold code in Graycode
The Gray code $G_{n}$ is divided into $l=2^{\xi}$ groups, where $\xi\leq n-1$. Each group consists of elements from $G_{n}$ that are spaced $l$ apart. Specifically, for each $l$ , the Gray code $G_{n}$ is partitioned such that the $i$-th group contains all elements of $G_{n}$ with indices satisfying $j\equiv i\mod l$, for $i=0,1,...,l-1$. Each group contains $2^{n-\xi}$ elements. We define the numbers within a group to form a Gray code shape if, aside from $\xi$ fixed bit positions, the remaining bits constitute a Gray code. The values of the fixed bits can either remain constant or undergo a single bit flip (0 $\leftrightarrow$ 1) between any two consecutive numbers in the sequence. We will prove by induction that each such group forms a Gray code shape. Consequently, this demonstrates that $G_{n}$ is suitable for encoding the operator $b^{l}$.
For $n=1$, the Gray code $G_{1}=(0,1).$ Dividing it into $l=2^{\xi}$ groups with $\xi=0,$ we get a single group containing $22^{1}=$ 2 elements. Clearly, each group trivially forms a Gray code shape. Assume that for some $n=k$, the Gray code $G_{k}$ can be partitioned into $l=2^{\xi}$groups, where each group forms a Gray code shape. Now, consider $n=k+1.$ The Gray code $G_{k+1}$ is constructed using the recursive formula $G_{k+1}=(\boldsymbol{0}\cdot G_{k},\boldsymbol{1}\cdot\bar{G}{}_{k})$. By the inductive hypothesis, we know that $G_{k}$ is partitioned into $l=2^{\xi}$ groups, and each group forms a Gray code shape. When we extend to $G_{k+1}$, the same partitioning scheme applies. Each of the $2^{\xi}$ groups in $G_{k}$ gives rise to two corresponding groups in $G_{k+1}$ by by appending 00 or $1$ to the front of the elements in $G_{k}$. The bit flip between the two groups in $G_{k+1}$ only affects the leading bit, leaving the remaining bits from $G_{k}$ unaffected. Therefore, each of the $l$ groups is extended by directly connecting the corresponding group in $\boldsymbol{0}\cdot G_{k}$ and $\boldsymbol{1}\cdot\bar{G}{}_{k}$.
  • Base Case ($n=1$)
    For $n=1$, the Gray code is given by: $G_{1}=(0,1)$. Here, we must have $\xi=0$ (since $\xi\leq n-1$), so $l=2^{0}=1$. Thus, the entire Gray code forms a single group containing $22^{1}=$ 2 elements. In this case, the group trivially forms a Gray code shape since the two elements differ in exactly one bit.
  • Inductive Hypothesis
    Assume that for some $n=k$, the Gray code $G_{k}$ can be partitioned into $l=2^{\xi}$ groups such that in each group, aside from $\xi$ fixed bits, the remaining $k-\xi$ bits form a Gray code. That is, in every group, consecutive elements differ in exactly one bit in the non-fixed positions, and any change in the fixed positions occurs as a single bit flip.
  • Inductive Step ($n=k+1$)
    The $(k+1)$-bit Gray code is constructed using the recursive formula: $G_{k+1}=(\boldsymbol{0}\cdot G_{k},\boldsymbol{1}\cdot\bar{G}{}_{k})$. We partition $G_{k+1}$ into $l=2^{\xi}$ groups by taking elements with indices congruent modulo $l$. Notice that:
    1. 1.
      Within the First Half $(\boldsymbol{0}\cdot G_{k})$: Each group in $G_{k}$(by the inductive hypothesis) forms a Gray code shape. Prefixing these elements with 00 merely adds a fixed bit to the beginning of each element without affecting the Gray code property in the remaining $k-\xi$ bits.
    2. 2.
      Within the Second Half $(\boldsymbol{1}\cdot\bar{G}{}_{k})$: Although this half is constructed from the reversed $G_{k}$, the reversal does not alter the property that consecutive elements differ in one bit in the non-fixed positions. Prefixing these elements with 1 similarly introduces a fixed bit that changes only in a controlled manner (at most once between consecutive elements).
    3. 3.
      At the Junction Between the Two Halves: Consider the transition from the last element of $\boldsymbol{0}\cdot G_{k}$ to the first element of $\boldsymbol{1}\cdot\bar{G}{}_{k}$ within the same group. By the standard construction of Gray codes, this transition involves a flip in the leading bit only, while the remaining $k$ bits (inherited from $G_{k}$ or its reverse) still satisfy the Gray code property (i.e., they differ in exactly one bit as dictated by the inductive hypothesis).
Thus, in $G_{k+1}$, each of the $2^{\xi}$ groups is obtained by “extending” the corresponding group from $G_{k}$ in both halves (with the 0-prefix and 1-prefix, respectively). The only additional difference between the two halves is in the newly added leading bit, and its change is isolated to a single bit flip. Therefore, aside from the $\xi$ fixed bits (now including the additional fixed prefix in each half), the remaining bits in each group form a Gray code sequence.
Appendix C Circuit Transpilation Example
[…]
Figure 8: The original circuit designed to compute the trace $\text{tr}\left(Y_{1}\otimes X_{0}\right)$. The circuit includes the ansatz $\Gamma(\boldsymbol{\theta})$ with the specified parameters for $t=0.5.$
To illustrate the transpilation process, we provide an explicit example for the circuit calculating $\text{tr}\left(Y_{1}\otimes X_{0}\right)$. The original circuit (Fig. 8) includes a parameterized ansatz $\Gamma(\boldsymbol{\theta})=\Gamma(\theta_{0},\theta_{1},\theta_{2},\theta_{3})$ followed by gates for Pauli measurement. The corresponding parameters shown in Fig. 8 for $t=0.5$. While functionally correct, this circuit uses generic gates unsuited for direct execution on the Zuchongzhi-2 platform.
[…]
Figure 9: The transpiled version of the circuit for the quantum processor hardware. The circuit is now in a format compatible with the basic gate set, ensuring that no further decomposition is required on the cloud server.
Using the bqskit toolkit, we transpile the circuit into a hardware-optimized form (Fig. 9). This process maps all operations to the device’s native gate set $\{X2P,X2M,Y2P,Y2M,R_{z},CZ\}$ while preserving the logical functionality. Though the transpiled circuit appears longer, it avoids further decomposition on the cloud server of the quantum hardware, reducing error accumulation. The toolkit occasionally generates different transpiled circuits, depending on the optimization process. Researchers may further optimize gate sequences using toolkit’s advanced compilation strategies, though the provided version already ensures reliable execution.